Suppose the own price elasticity of demand for good X is 2, its income elasticity is 3, its advertising elasticity is 4, and the cross-price elasticity of demand between it and good Y

is 6. Determine how much the consumption of this good will change if: (a) The price of good X increased by 5 percent, (b) the price of good Y increase by 10 percent, (c) Advertising decreases by 2 present, (d) income falls by 3 percent.



1
Expert's answer
2021-07-20T15:27:06-0400

a.

The price of good X decreases by 5 percent.

"Own\space price\space elasticity \space of\space demand =\frac{ percentage \space change \space in \space demand }{ percentage \space change \space in\space price.}"

"2 = \frac{percentage \space change\space in\space demand }{ -5\%}"

-10% = percentage change in demand (A decrease).


b.

The price of good Y increases by 10 percent.

"Cross\space price \space elasticity \space of\space demand =\frac{ percentage \space change \space in\space demand \space of\space good \space X }{ percentage \space change\space in\space price\space of\space good \space Y}."

"6 =\frac{ percentage\space change\space in\space demand\space of\space good \space X }{ 8\%}"

48% = percentage change in demand of good X (An increase).


c.

 Advertising decreases by 2 percent.

"Advertising\space elasticity\space of\space demand = \frac{percentage\space change\space in\space demand\space of\space good\space X }{percentage\space change\space in \space advertising}"

"4 = \frac{percentage change in demand of good X }{ -4\%}"

-16% = percentage change in demand (A decrease).


d.

Income decreases by 3 percent.

"Income\space elasticity \space of\space demand =\frac{ percentage \space change \space in \space demand\space of\space good \space X }{ percentage\space change\space in\space income}"

"3 =\frac{ percentage\space change\space in\space demand \space of\space good\space X }{ -3\%}"

-9% = percentage change in demand of good X (A decrease)